# Symmetry-informed transferability of optimal parameters in the Quantum Approximate Optimization Algorithm

### v1

@@ IV Conclusions and outlook
In this article, we extend the known energy landscape symmetries of MaxCut and show how to leverage them to transfer optimal QAOA parameters. While previous studies focus on unweighted graphs, we find that for all graphs where the sum of the weights connected to each vertex is even, the landscape follows the EWS symmetry (see Table 1). Analogously, the landscape obeys the OWS symmetry for all graphs where the sum of the weights connected to each vertex is odd. Among the examples following these symmetries, we find regular graphs with weights $w_{ij}\in\{-1,1\}$.
@@
Regarding the optimal parameters for $p=1$, it is known that unweighted triangle-free regular graphs have an optimal set in the domain $\mathcal{U}_{1}=[0,\pi/2)\times[0,\pi/4)$ [43], and from the symmetries, other optimal sets in the green solid areas of Fig. 1, depending on the parity of the degree. Here, we analyze the transferability of these optimal parameters, i.e. those of a $d^{\prime}$-regular tree subgraph, to triangle-free $d$-regular graphs, $d$-regular graphs, different kinds of random graphs, and $d$-regular graphs with integer weights. Our work shows that, in all cases, the graphs are viable recipients for the transferred parameters.
@@
In future work, our study on symmetries and transferability can be extended to higher algorithm depths and other kinds of problems and versions of VQAs, such as the multiangle QAOA, with different circuit parameterizations. These efforts may lead to further simplifications of the classical optimization task in VQAs, one of the most challenging aspects of their successful implementation.
@@ Acknowledgements.
We use the Python-based packages QTensor [49] and Qiskit [50] for the numerical simulations of QAOA and Matplotlib [51] for data visualization. We acknowledge useful discussions with Zeidan Zeidan and Göran Johansson. We acknowledge support from the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). L.G.-Á. further acknowledges funding under Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA via the project 101113946 OpenSuperQPlus100.
Appendix A Symmetries of optimal QAOA parameters on MaxCut
@@
A.1 Symmetries for graphs where the sum of weights connected to each vertex is even
In this section, we prove the symmetry called EWS in Section II. We consider a variational state constructed by a QAOA circuit of depth $p$. We focus on the $m$-th iteration of the algorithm, in which the operation $\hat{U}_{M}(\beta_{m})\hat{U}_{C}(\gamma_{m})$ acts on the intermediate variational state
@@ A.1 Symmetries for graphs where the sum of weights connected to each vertex is even
A.2 Symmetries for graphs where the sum of weights connected to each vertex is odd
In this section, we prove the symmetry called OWS in Section II. As in the previous case, we consider the general case of a QAOA circuit of depth $p$ and how the change $\gamma_{m}\rightarrow\gamma_{m}\pm\pi$ in the $m$-th layer changes the final output state. Given the cost Hamiltonian $\hat{H}_{C}$ of Eq. (3) for MaxCut on a weighted graph, the algorithm operation of layer $m$ is
@@ A.2 Symmetries for graphs where the sum of weights connected to each vertex is odd
We only consider graphs with integer weights $w_{ij}$, which can be either even or odd. If a given weight $w_{ij}$ is even, only the cosine differs from zero, which just adds a global phase. Otherwise, if a weight $w_{ij}$ is odd, only the sine function is different from zero, and the resulting operation of the $m$-th layer in QAOA changes. For convenience, we rewrite Eq. (A.2) as
@@
$e^{i\kappa}$ $=\prod_{\mathclap{(i,j)\in E}}\exp(\mp i\frac{w_{ij}\pi}{2})\prod_{\mathclap{\begin{subarray}{c}(k,\ell)\in E\\ w_{k\ell}\textrm{ even}\end{subarray}}}\cos(\frac{w_{k\ell}\pi}{2}),$
$e^{i\delta}$ $=\mp i\prod_{\mathclap{\begin{subarray}{c}(i,j)\in E\\ w_{ij}\textrm{ odd}\end{subarray}}}\sin(\frac{w_{ij}\pi}{2}).$
@@
and we have two cases: if $\mu_{k}$ is even, then $(\hat{\sigma}_{z}^{k})^{\mu_{k}}=\mathbb{1}$; otherwise, if $\mu_{k}$ is odd, then $(\hat{\sigma}_{z}^{k})^{\mu_{k}}=\hat{\sigma}_{z}^{k}$. In the first case, with $\mu_{k}$ even for all $k\in V$, we recover again the symmetry rule $\gamma_{m}\rightarrow\gamma_{m}\pm\pi$ described in the previous section, since the sum of all weights connected to each vertex is even.
@@
Using the relation $\hat{\sigma}_{z}^{k}\hat{\sigma}_{x}^{k}=-\hat{\sigma}_{x}^{k}\hat{\sigma}_{z}^{k}$, and $e^{i\theta}=e^{i\kappa}e^{i\delta}$, we can rewrite Eq. (A.2) as
@@
with $\bm{\gamma}^{\prime}=(\gamma_{1},\dots,\gamma_{m-1},\gamma_{m}\pm\pi,\gamma_{m+1},\dots,\gamma_{p})$ and $\bm{\beta}^{\prime}=(\beta_{1},\dots,\beta_{m-1},\beta^{\prime}_{m},\beta^{\prime}_{m+1},\dots,\beta^{\prime}_{p})$. Using again the commutation relations of the Pauli operators, we can rewrite Eq. (A.2) as
@@
Thus, graphs where every vertex has an odd number of odd weighted edges connected to it are symmetric under the parameter changes of Eq. (A.2). In other words, graphs where the sum of weights connected to each vertex is odd will exhibit these parameter symmetries.
@@
Appendix B Derivation of the analytic expression of the transferability error

### v2

@@ V Conclusions and outlook
Optimizing the variational circuits remains a big obstacle for realizable and successful VQAs. One useful simplification in the optimization task is limiting the search space using symmetries in the energy landscape. Additionally, the optimal parameters concentrate on typical values for different instances of certain problems. Thus, we can use these typical values as pretrained parameters and transfer them between instances, either to be used directly, or as initialization points for a warm-start optimization. In this article, we extend the known energy landscape symmetries of MaxCut and show how to leverage them to transfer optimal QAOA parameters. While previous studies focus on unweighted graphs, we find that for all graphs where the sum of the weights connected to each vertex is even, the landscape follows the EWS symmetry (see Table 1). Analogously, the landscape obeys the OWS symmetry for all graphs where the sum of the weights connected to each vertex is odd. Among the examples following these symmetries, we find regular graphs with weights $w_{ij}\in\{-1,1\}$.
@@
Regarding the optimal parameters for $p=1$, it is known that unweighted triangle-free regular graphs have an optimal set in the domain $\mathcal{U}_{1}=[0,\pi/2)\times[0,\pi/4)$ [52], and from the symmetries, other optimal sets in the green solid areas of Fig. 1, depending on the parity of the degree. Here, we analyze the transferability of these optimal parameters, i.e., those of a $d^{\prime}$-regular tree subgraph, to triangle-free $d$-regular graphs, $d$-regular graphs, different kinds of random graphs, and $d$-regular graphs with integer weights. Our work shows that, in all cases, the graphs are viable recipients for the transferred parameters.
@@
Finally, we discuss the applicability of our results to other COPs and VQAs. We find conditions such that adding a local field term to the cost Hamiltonian in QAOA preserves the two symmetries EWS and OWS, thus including general COPs described by Ising Hamiltonians [58]. Additionally, we show that the EWS is valid for all unitaries associated with Hamiltonians with $\hat{\sigma}_{x}\hat{\sigma}_{x}$, $\hat{\sigma}_{y}\hat{\sigma}_{y}$, or $\hat{\sigma}_{z}\hat{\sigma}_{z}$ interactions. Among these Hamiltonians, we consider the general case of the Heisenberg model on high-dimensional spaces with integer-valued couplings. For specific coupling values, this model reduces to the XYZ and XXZ models, previously studied with glshva [57]. We also examine two QAOA variants, RQAOA and ma-QAOA, which improve the performance of the standard QAOA but require more parameter optimizations. We find that the iterative process of RQAOA only maintains the EWS, but breaks the OWS after the first iteration. Furthermore, we observe that the ma-QAOA could benefit from specific transferability schemes based on the symmetries of individual subgraphs, instead of the entire graph.
Our insights on the conditions of transferred parameters and their relation to the problem symmetries set the stage for a more comprehensive study of symmetry-based optimization in quantum algorithms. Future work may seek to numerically investigate the benefits of reducing the landscape size with these symmetries and performing warm-start local optimization with transferred parameters. For instance, the next steps would involve characterizing the number of iterations in the classical optimization compared to the current standard approaches.
@@ Acknowledgements.
We acknowledge useful discussions with Zeidan Zeidan and Göran Johansson of QAOA and our results. We acknowledge useful discussions with Werner Dobrautz and Erika Magnusson of the VQAs used in chemistry. We acknowledge support from the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). L.G.-Á. further acknowledges funding under Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA via the project 101113946 OpenSuperQPlus100.
Appendix A Simulation methods
In this section, we detail the methods used to simulate the QAOA and perform the classical optimization. All graphs are generated using the Python package NetworkX [61], and the figures are created with the Python package Matplotlib [62]. We simulate MaxCut with $d$-regular graphs and random graphs on QAOA using the tensor-network-based package Qtensor [63, 64]. The optimization of the parameters was done with the Qtensor implementation of the Python package PyTorchs [65] optimizer RMSprop, with 50 optimization steps.
We simulated MaxCut on weighted $d$-regular graphs with QAOA using Qiskit [66], with 3000 shots per circuit evaluation. Here, the parameters’ optimization was carried out with the Python package SciPys [67] and the optimizer differential_evolution with a population size of 15, and a maximum of 1000 function calls. For all optimizations, we used the optimal parameters to the relevant $d$-regular tree subgraphs as initial guesses.
Appendix B Symmetries of optimal QAOA parameters on MaxCut
@@
B.1 Symmetries for graphs where the sum of weights connected to each vertex is even
In this section, we prove the symmetry called EWS in Sec. II. We consider a variational state constructed by a QAOA circuit of depth $p$. We focus on the $m$-th iteration of the algorithm, in which the operation $\hat{U}_{M}(\beta_{m})\hat{U}_{C}(\gamma_{m})$ acts on the intermediate variational state
@@ B.1 Symmetries for graphs where the sum of weights connected to each vertex is even
B.2 Symmetries for graphs where the sum of weights connected to each vertex is odd
In this section, we prove the symmetry called OWS in Sec. II. As in the previous case, we consider the general case of a QAOA circuit of depth $p$ and how the change $\gamma_{m}\rightarrow\gamma_{m}\pm\pi$ in the $m$-th layer changes the final output state. Given the cost Hamiltonian $\hat{H}_{C}$ of Eq. (3) for MaxCut on a weighted graph, the algorithm operation of layer $m$ is
@@ B.2 Symmetries for graphs where the sum of weights connected to each vertex is odd
We only consider graphs with integer weights $w_{ij}$, which can be either even or odd. If a given weight $w_{ij}$ is even, only the cosine differs from zero, which just adds a global phase. Otherwise, if a weight $w_{ij}$ is odd, only the sine function is different from zero, and the resulting operation of the $m$-th layer in QAOA changes. For convenience, we rewrite Eq. (B.2) as
@@
$e^{i\kappa}$ $=\prod_{\mathclap{(i,j)\in E}}\exp(\mp i\frac{w_{ij}\pi}{2})\prod_{\mathclap{\begin{subarray}{c}(k,\ell)\in E\\ w_{\ell k}\textrm{ even}\end{subarray}}}\cos(\frac{w_{\ell k}\pi}{2}),$ (18)
$e^{i\delta}$ $=\prod_{\mathclap{\begin{subarray}{c}(i,j)\in E\\ w_{ij}\textrm{ odd}\end{subarray}}}i\sin(\mp\frac{w_{ij}\pi}{2}).$ (19)
@@
and we have two cases: if $\mu_{k}$ is even, then $(\hat{\sigma}_{z}^{k})^{\mu_{k}}=\mathbb{1}$; otherwise, if $\mu_{k}$ is odd, then $(\hat{\sigma}_{z}^{k})^{\mu_{k}}=\hat{\sigma}_{z}^{k}$. In the first case, with $\mu_{k}$ even for all $k\in V$, we recover again the symmetry rule $\gamma_{m}\rightarrow\gamma_{m}\pm\pi$ described in the previous section, since the sum of all weights connected to each vertex is even. We note that if the cost Hamiltonian was expressed with $\hat{\sigma}_{x}\hat{\sigma}_{x}$ or $\hat{\sigma}_{y}\hat{\sigma}_{y}$ products instead of $\hat{\sigma}_{z}\hat{\sigma}_{z}$, similar derivations would hold. We discuss this fact further in our analyses of other VQAs in Appendix D.
@@
Using the relation $\hat{\sigma}_{z}^{k}\hat{\sigma}_{x}^{k}=-\hat{\sigma}_{x}^{k}\hat{\sigma}_{z}^{k}$, and $e^{i\theta}=e^{i\kappa}e^{i\delta}$, we can rewrite Eq. (B.2) as
@@
with $\bm{\gamma}^{\prime}=(\gamma_{1},\dots,\gamma_{m-1},\gamma_{m}\pm\pi,\gamma_{m+1},\dots,\gamma_{p})$ and $\bm{\beta}^{\prime}=(\beta_{1},\dots,\beta_{m-1},\beta^{\prime}_{m},\beta^{\prime}_{m+1},\dots,\beta^{\prime}_{p})$. Using again the commutation relations of the Pauli operators, we can rewrite Eq. (B.2) as
@@
Thus, graphs where every vertex has an odd number of odd weighted edges connected to it are symmetric under the parameter changes of Eq. (B.2). In other words, graphs where the sum of weights connected to each vertex is odd will exhibit these parameter symmetries.
@@
Appendix C Symmetries of optimal QAOA parameters on general COPs
In this section, we extend the results of Appendix B and find the conditions for which the symmetries EWS and OWS, introduced in Sec. II, still apply if the cost Hamiltonian includes local field terms. In other words, a general cost Hamiltonian in the form of an Ising Hamiltonian can result in the parameter landscape having EWS or OWS. Many COPs can be mapped to an Ising Hamiltonian, including all of Karp’s 21 NP-complete problems [58].
A general Ising Hamiltonian with local field terms reads
$\hat{H}_{C}=\frac{1}{2}\sum_{\mathclap{(i,j)\in E}}w_{ij}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{j}+\frac{1}{2}\sum_{i\in V}h_{i}\hat{\sigma}_{z}^{i},$ (29)
where $E$ is the set of edges in an interaction graph, and $V$ is the set of spins. The QAOA unitary operation associated to this cost Hamiltonian is
$\hat{U}_{C}(\gamma_{m})$ $=\prod_{\mathclap{(i,j)\in E}}\exp(-i\frac{w_{ij}}{2}\gamma_{m}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{j})$
×$\times\prod_{\mathclap{i\in V}}\exp(-i\frac{h_{i}}{2}\gamma_{m}\hat{\sigma}_{z}^{i}).$ (30)
We now consider the parameter shift $\gamma_{m}\pm\pi$, so that
$\hat{U}_{C}(\gamma\pm\pi)$ $=\hat{U}_{C}(\gamma_{m})\prod_{\mathclap{(i,j)\in E}}\exp(\mp i\frac{w_{ij}\pi}{2}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{j})$
×$\times\prod_{\mathclap{i\in V}}\exp(\mp\frac{ih_{i}\pi}{2}\hat{\sigma}_{z}^{i}).$ (31)
Analogously to Appendix B, we assume that the coefficients $w_{ij}$ and $h_{i}$ are integers. Then, we can rewrite the terms of Eq. (C) as
$\prod_{\mathclap{(i,j)\in E}}\exp(\mp i\frac{w_{ij}\pi}{2}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{j})=e^{i\varepsilon}e^{i\delta}\prod_{k\in V}(\hat{\sigma}_{z}^{k})^{\mu_{k}},$ (32)
$\exp(\mp\frac{ih_{i}\pi}{2}\hat{\sigma}_{z}^{i})=\cos\left(\frac{h_{i}\pi}{2}\right)\mp i\sin\left(\frac{h_{i}\pi}{2}\right)\hat{\sigma}_{z}^{i},$ (33)
where
$e^{i\varepsilon}=\prod_{\mathclap{\begin{subarray}{c}(k,\ell)\in E\\ w_{\ell k}\textrm{ even}\end{subarray}}}\cos(\frac{w_{\ell k}\pi}{2}),$ (34)
and $e^{i\delta}$ is given by Eq. (19). For integer coefficients $h_{i}$, the right-hand side of Eq. (33) is $\mathbb{1}$ if $h_{i}$ is even, and $\mp i\hat{\sigma}_{z}^{i}$ if $h_{i}$ is odd. Thus, we get
$\prod_{\mathclap{i\in V}}\exp(\mp\frac{ih_{i}\pi}{2}\hat{\sigma}_{z}^{i})=e^{i\lambda}e^{i\nu}\prod_{\mathclap{k\in V}}(\hat{\sigma}_{z}^{k})^{h_{k}},$ (35)
with the global phases
$e^{i\lambda}=\prod_{\mathclap{\begin{subarray}{c}i\in V\\ h_{i}\textrm{ odd}\end{subarray}}}i\sin(\mp\frac{h_{i}\pi}{2}),\textrm{ and }e^{i\nu}=\prod_{\mathclap{\begin{subarray}{c}i\in V\\ h_{i}\textrm{ odd}\end{subarray}}}\cos(\frac{h_{i}\pi}{2}).$ (36)
We can finally rewrite Eq. (C) as
$\hat{U}_{C}(\gamma_{m}\pm\pi)=e^{i\alpha}\hat{U}_{C}(\gamma_{m})\prod_{\mathclap{k\in V}}(\hat{\sigma}_{z}^{k})^{\mu_{k}+h_{k}},$ (37)
with the global phase $e^{i\alpha}=e^{i\varepsilon}e^{i\delta}e^{i\lambda}e^{i\nu}$. Now, we can follow a similar argument as in Appendix B considering the quantity $\mu_{k}+h_{k}$, instead of just $\mu_{k}$, which was the number of odd-valued weights connected to vertex $k$. We conclude that the EWS applies if $\mu_{k}+h_{k}$ is even-valued, and OWS applies if $\mu_{k}+h_{k}$ is odd-valued.
Appendix D Symmetries of optimal parameters in other VQAs
Here, we connect and extend the symmetries derived for the QAOA for the MaxCut problem to the HVA for condensed matter problems, such as those described by Eq. (9) in Sec. IV. Moreover, we include a detailed derivation for similar symmetry results in iterative quantum algorithms such as the RQAOA.
D.1 Symmetries for interaction graphs where the sum of weights connected to each vertex is even
We consider the unitary operator from Eq. (8),
$\hat{U}_{\alpha\alpha}(\gamma)=\prod_{\mathclap{(i,j)\in E}}\exp(-i\frac{w_{ij}}{2}\hat{\sigma}_{\alpha}^{i}\hat{\sigma}_{\alpha}^{j}),$ (38)
with $\alpha=\{x,y,z\}$, and $E$ the set of edges in an interaction graph. We assume that the weights satisfy $\sum_{j}w_{jk}=2\mathbb{Z},\forall k\in V$, where $V$ is the set of all spins with an interaction. Now, following the same line of arguments made in Appendix B, we find that
$\hat{U}_{\alpha\alpha}(\gamma\pm\pi)=$ $\prod_{\mathclap{(i,j)\in E}}\exp(-i\frac{w_{ij}}{2}\hat{\sigma}_{\alpha}^{i}\hat{\sigma}_{\alpha}^{j})$
$=$ $\,\hat{U}_{\alpha\alpha}(\gamma)\prod_{\mathclap{(i,j)\in E}}\exp(\mp i\pi\frac{w_{ij}}{2}\hat{\sigma}_{\alpha}^{i}\hat{\sigma}_{\alpha}^{j})$
$=$ $\,\hat{U}_{\alpha\alpha}(\gamma)e^{i\kappa}e^{i\delta}\prod_{\mathclap{\begin{subarray}{c}(i,j)\in E\\ w_{ij}\textrm{ odd}\end{subarray}}}\hat{\sigma}_{\alpha}^{i}\hat{\sigma}_{\alpha}^{j}$
$=$ $\,\hat{U}_{\alpha\alpha}(\gamma)e^{i\kappa}e^{i\delta}\prod_{\mathclap{k\in V}}(\hat{\sigma}_{\alpha}^{k})^{\mu_{k}}.$ (39)
The global phases $e^{i\kappa}$ and $e^{i\delta}$ are given by Eq. (18) and (19) respectivly, $\mu_{k}$ is the number of odd-valued weights of the set $E_{k}$. The initial assumption made on the weights imply that $\mu_{k}$ is even, so $(\hat{\sigma}_{\alpha}^{k})^{\mu_{k}}=\mathbb{1}$, and we retrieve that, up to a global phase,
$\hat{U}_{\alpha\alpha}(\gamma\pm\pi)=\hat{U}_{\alpha\alpha}(\gamma).$ (40)
It follows that ansätze made up of products of operators on the form of Eq. 38 will have EWS, like the one-dimensional Heisenberg XYZ model.
D.2 Symmetries for RQAOA
In this section, we show that RQAOA preserves EWS, but breaks OWS. The RQAOA is a recursive algorithm in which, after each iteration, one of the vertices with the largest correlation magnitude is effectively removed from the graph. Formally, if the maximal correlation is $|\mathcal{M}_{k\ell}|$, with $\mathcal{M}_{k\ell}=\bra{\bm{\gamma},\bm{\beta}}\hat{\sigma}_{z}^{k}\hat{\sigma}_{z}^{\ell}\ket{\bm{\gamma},\bm{\beta}}$, one of the vertices $k$ or $\ell$ is removed by imposing the constraint $\hat{\sigma}_{z}^{\ell}=\text{sgn}(M_{k\ell})\hat{\sigma}_{z}^{k}$. For our purpose, it is sufficient to consider $\hat{\sigma}_{z}^{\ell}=\pm\hat{\sigma}_{z}^{k}$.
With RQAOA, the dimension of the Hilbert space decreases every iteration. Thus, we can identify each distinct cost Hamiltonian of different iterations by the systems’ size or number of qubits, $n$. Here, we label the initial Hamiltonian as $\hat{H}_{n}$, the Hamiltonian after the recursive step, $\hat{H}_{n-1}$, and so on. We are interested in how the sum of integer coefficients (weights and local field) connected to each vertex changes with each iteration. We recall from Sec. IV that, for graphs with EWS, the sum of coefficients connected to any vertex $k$ is even, $\Sigma_{j}w_{jk}+h_{k}\in 2\mathbb{Z},\forall k\in V$. In contrast, for graphs with OWS, the sum of coefficients connected to any vertex $k$ is odd, $\Sigma_{j}w_{jk}+h_{k}\in\mathbb{Z}\setminus 2\mathbb{Z},\forall k\in V$.
Now, we consider an initial cost Hamiltonian $\hat{H}_{n}$ related to a graph that obeys either EWS or OWS. After the first iteration, we impose the constraint $\hat{\sigma}_{z}^{\ell}=\pm\hat{\sigma}_{z}^{k}$ such that the vertex $\ell$ is removed from the graph—or qubit $\ell$ is removed from the system. To analyze the symmetry of the new Hamiltonian $\hat{H}_{n-1}$, we first examine the terms of the initial cost Hamiltonian related to each vertex $i$, that is, the terms in which qubit $i$ participates. We denote such terms as $\hat{H}_{n}^{i}=\sum_{j}w_{ij}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{j}+h_{i}$, and the total cost Hamiltonian can be written then as $\hat{H}_{n}=\sum_{i}\hat{H}_{n}^{i}/2$.
For a given vertex $i$, we find that there are three possibilities: either $i$ share no edge with the removed vertex $\ell$, or $i$ share one edge with the removed vertex $\ell$, or $i=k$, meaning that it shares one edge with the removed vertex $\ell$, and that edge is removed by the constraint. In the first case, if the vertex $i$ does not share an edge with vertex $\ell$, we observe that $\hat{H}_{n-1}^{i}=\hat{H}_{n}^{i}$. Second, if the vertex $i$ shares an edge with vertex $\ell$ with the weight $w_{i\ell}$, after the RQAOA iteration, this vertex $i$ will share a new edge with vertex $k$ instead with weight ±$\pm w_{i\ell}$. Then, the initial Hamiltonian terms
$\hat{H}_{n}^{i}=\sum_{j\neq\ell}w_{ij}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{j}+w_{i\ell}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{\ell}+h_{i}\hat{\sigma}_{z}^{i},$ (41)
will change to
$\hat{H}_{n-1}^{i}=\sum_{j\neq\ell}w_{ij}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{j}\pm w_{i\ell}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{k}+h_{i}\hat{\sigma}_{z}^{i}.$ (42)
Let us consider the difference in the sum of coefficients between Eq. (41) and (42). The weights of the edges not connected to vertex $\ell$ and the local field $h_{i}$ remain unchanged, so the difference is given by $w_{i\ell}\pm w_{i\ell}$. This quantity is always even, as it is either equal to zero, or to $2w_{i\ell}$. Thus, the parity of the sum of coefficients around a vertex $i$ in this second case does not change by the constraint imposed in the recursive algorithm. Lastly, we consider the vertices directly subject to the constraint, $\ell$ and $k$. The contribution to the initial cost Hamiltonian related to vertex $k$ is
$\hat{H}_{n}^{k}=\sum_{j\neq\ell}w_{jk}\hat{\sigma}_{z}^{j}\hat{\sigma}_{z}^{k}+w_{\ell k}\hat{\sigma}_{z}^{\ell}\hat{\sigma}_{z}^{k}+h_{k}\hat{\sigma}_{z}^{k},$ (43)
while for vertex $\ell$ we have
$\hat{H}_{n}^{\ell}=\sum_{j\neq k}w_{j\ell}\hat{\sigma}_{z}^{j}\hat{\sigma}_{z}^{\ell}+w_{\ell k}\hat{\sigma}_{z}^{k}\hat{\sigma}_{z}^{\ell}+h_{\ell}\hat{\sigma}_{z}^{\ell}.$ (44)
As the graph follow either the EWS or the OWS, the parity of the sums of coefficients in Eqs. (43) and (44) is the same. After the first RQAOA iteration, we get
$\hat{H}_{n-1}^{k}$ $=\sum_{j\neq\ell}w_{jk}\hat{\sigma}_{z}^{j}\hat{\sigma}_{z}^{k}\pm\sum_{j\neq k}w_{j\ell}\hat{\sigma}_{z}^{j}\hat{\sigma}_{z}^{k}+(w_{\ell k}\pm w_{\ell k})$
$+(h_{k}\pm h_{\ell})\hat{\sigma}_{z}^{k},$ (45)
and $\hat{H}_{n-1}^{\ell}=0$. The term $(w_{\ell k}\pm w_{\ell k})$ in Eq. (D.2) is a constant energy shift, and not a weight on an edge.
We focus now on the sum of coefficients connected to vertex $k$ in the Hamiltonian $\hat{H}_{n-1}^{k}$ of Eq. (D.2), denoted as $s_{n-1}(k)$, which is
$s_{n-1}(k)=$ $\sum_{j\neq\ell}w_{jk}\pm\sum_{j\neq k}w_{j\ell}+(h_{k}\pm h_{\ell})$
$=$ $\sum_{j\neq\ell}w_{jk}+w_{\ell k}+h_{k}\pm\sum_{j\neq k}w_{j\ell}\pm w_{\ell k}\pm h_{\ell}$
-- $w_{\ell k}\mp w_{\ell k},$ (46)
where we have added and subtracted $w_{\ell k}$ twice. The first three terms of the second line in Eq. (D.2) are equal to $s_{n}(k)$, the sum of coefficients that connected to vertex $k$ in $\hat{H}_{n}^{k}$. Similarly, the next three terms correspond to ±$\pm s_{n}(\ell)$. We get
$s_{n-1}(k)=s_{n}(k)\pm s_{n}(\ell)-w_{\ell k}\mp w_{\ell k}.$ (47)
As the graph in $\hat{H}_{n}$ is assumed to be either EWS or OWS, $s_{n}(k)$ and $s_{n}(\ell)$ have the same parity, so the sum, or difference, of these quantities must be even. The quantity $-w_{\ell k}\mp w_{\ell k}$ is also even, making Eq. (47) even.
To summarize, after the constraint is imposed, all vertices except one are related to coefficients that sum to a value with unchanged parity. The exception is vertex $k$, the one directly affected by the constraint in the previous example. Regardless of the symmetry of the initial Hamiltonian $\hat{H}_{n}$, the sum of coefficients associated with it after the iteration, $s_{n-1}(k)$, is even. As such, RQAOA preserves EWS, but breaks OWS.
Appendix E Derivation of the analytic expression of the transferability error